OpenStudy (anonymous):

A particle has a charge of 5.2..*.10ȹ© coulombs and experiences a force of 9.5..*.10ȹ¢ newtons when it travels through a magnetic field with strength 2.2..*.10ȹ tesla. What is the speed of the particle?

OpenStudy (irishboy123):

F = qv X B |F| = |qv| |B| sin ø assuming travels at right angles to the field (ø = π/2), "speed" = F/qB = (9.5 something,,,can't read it)/(5.2 ...something • 2.2,,,,something)

OpenStudy (anonymous):

ok